If the smallest of the three consecutive integer is added to twice the largest, the result is seven less than four times the middle. Find the nos
Answer by ankor@dixie-net.com(22740)

: If the smallest of the three consecutive integer is added to twice the largest, the result is seven less than four times the middle. Find the nos
:
The three consecutive integers: x, (x+1), (x+2)
:
x + 2(x+2) = 4(x+1) - 7
:
x + 2x + 4 = 4x + 4 - 7
:
3x + 4 = 4x - 3
:
3x - 4x = -3 - 4
:
-x = -7
:
x = +7, 8, 9 are the three integers
:
Check solution in the statement:
"If the smallest of the three consecutive integers, is added to twice the largest, the result is seven less than four times the middle. "
7 + 2(9) = 4(8) - 7
7 + 18 = 32 - 7; confirms our solution
:
Not that hard, right?